# DisTaC: Conditioning Task Vectors via Distillation for Robust Model Merging

### v1

@@ 8 Conclusion
We presented DisTaC, a lightweight and practical pre-conditioning method for task vectors that improves the robustness of model merging in multi-task learning. Our analysis identified two major failure modes—norm disparity and low source-model confidence—that frequently occur in real-world merging scenarios. DisTaC addresses both issues simultaneously via KD on unlabeled data, requiring only minimal computational cost and no access to task labels. Through extensive experiments, we demonstrated that DisTaC not only recovers performance degraded by task vector scaling, but also enhances confidence in the source models without sacrificing generalization. Furthermore, we showed that DisTaC enables state-of-the-art merging methods to succeed in challenging cases where they would otherwise fail. Our findings highlight the importance of task vector conditioning, and we believe that DisTaC provides a simple yet powerful tool to make model merging more reliable and broadly applicable.
Acknowledgments
This work was (partly) achieved through the use of large-scale computer systems at D3 Center, The University of Osaka. Hiroki Naganuma acknowledges funding support from the ANRI Fellowship for this work.
@@ Acknowledgments
Appendix A Theoretical Insights into Task Vector Merging for Models Optimized with Distinct Objectives
@@
A.1 Notation and Assumptions
We begin by formalizing the notation and assumptions used throughout this analysis.
Notation
  • Base parameter vector: $\bm{\theta}_{0}\in\mathbb{R}^{d}$. This is the shared starting point for fine-tuning, e.g., a pre-trained model’s parameters.
  • Tasks and datasets: Two distinct downstream tasks, $T_{1}$ and $T_{2}$, with corresponding datasets $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$.
  • Cross-Entropy (CE) loss: The standard objective for task $i$ is defined as
    $J_{i}^{\mathrm{CE}}(\bm{\theta}):=-\,\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\bigl{[}\log p_{\bm{\theta}}(y\mid x)\bigr{]}.$
  • Calibrated objective: A modified objective function for task $i$ that includes a generic penalty term $\mathcal{C}_{i}(\bm{\theta})$  [^1] For example, a detailed description of evaluating focal loss can be found in Kimura & Naganuma (2025), weighted by $\lambda_{i}>0$.
@@ A.1 Notation and Assumptions
  • Task-Specific optima: The parameter vector that minimizes a given objective for task $i$. We denote this by
    $\bm{\theta}_{i}^{\star}:=\arg\min_{\bm{\theta}}J_{i}^{\star}(\bm{\theta}),\quad\text{where }\star\in\{\mathrm{CE,CAL}\}.$
  • Task vectors: The displacement from the base parameters $\bm{\theta}_{0}$ to the task-specific optimum $\bm{\theta}_{i}^{\star}$.
    $\bm{\tau}_{i}^{\star}:=\bm{\theta}_{i}^{\star}-\bm{\theta}_{0}.$
Assumptions
  1. 1.
    The objective functions $J_{i}^{\mathrm{CE}}$ and $J_{i}^{\mathrm{CAL}}$ are twice continuously differentiable ($C^{2}$) in a neighborhood of the base parameter vector $\bm{\theta}_{0}$.
  2. 2.
    The Hessian matrix of the CE loss evaluated at $\bm{\theta}_{0}$, denoted $H_{i}:=\nabla^{2}J_{i}^{\mathrm{CE}}(\bm{\theta}_{0})$, is positive-definite. This ensures that $\bm{\theta}_{0}$ is located in a convex region of the CE loss landscape.
  3. 3.
    We define the following gradients at the base point $\bm{\theta}_{0}$ for notational convenience:
    $\textbf{g}_{i}:=\nabla J_{i}^{\mathrm{CE}}(\bm{\theta}_{0}),\qquad\textbf{b}_{i}:=\nabla\mathcal{C}_{i}(\bm{\theta}_{0}).$
A.2 Derivation of the Standard Task Vector
@@
A.3 Derivation of the Calibrated Task Vector
@@
A.3.1 Gradient and Hessian at the Base Point
@@
A.3.2 Neumann Series Expansion of $\tilde{H}_{i}^{-1}$
@@
The inverse is then given by $\tilde{H}_{i}^{-1}=(I+\lambda_{i}H_{i}^{-1}A_{i})^{-1}H_{i}^{-1}$. We can expand the term $(I+\lambda_{i}H_{i}^{-1}A_{i})^{-1}$ using a Neumann series (Horn & Johnson, 2012), $(I+X)^{-1}=\sum_{k=0}^{\infty}(-X)^{k}$, which converges if the spectral radius of $X$ is less than 1. Assuming $\lambda_{i}$ is sufficiently small such that $\|\lambda_{i}H_{i}^{-1}A_{i}\|<1$, we have:
@@ A.3.3 Solving for $\tau_{i}^{\mathrm{CAL}}$
In many practical scenarios, especially after extensive pre-training, the initial gradient norm $\|\textbf{g}_{i}\|$ is small. Consequently, the term $\lambda_{i}H_{i}^{-1}A_{i}H_{i}^{-1}\textbf{g}_{i}$, which is of order $\mathcal{O}(\lambda_{i}\|\textbf{g}_{i}\|)$, is often negligible compared to the term $-\lambda_{i}H_{i}^{-1}\textbf{b}_{i}$, which is $\mathcal{O}(\lambda_{i})$. Under this simplifying assumption, we can define the first-order correction due to calibration as:
@@
A.4 Task Vector Merging
@@
A.5 Taylor Expansion of the CE Loss for Merged Vectors
@@ A.5.1 Merging of CE Vectors
The term for task $i$ itself ($i=1$ and analyzing $\bm{\tau}_{1}^{\mathrm{CE}}$, or $i=2$ and analyzing $\bm{\tau}_{2}^{\mathrm{CE}}$) is $-\alpha(\bm{\tau}_{i}^{\mathrm{CE}})^{\top}H_{i}\bm{\tau}_{i}^{\mathrm{CE}}=-\alpha\|\bm{\tau}_{i}^{\mathrm{CE}}\|_{H_{i}}^{2}$. Since $H_{i}$ is positive-definite, this self-term is strictly negative. The cross-term’s sign is indefinite. However, the dominant contribution to the loss change is typically negative and of order $\mathcal{O}(\|\bm{\tau}\|^{2})$, indicating that merging CE vectors does not increase the loss at first order.
A.5.2 Merging of Calibrated Vectors
@@
A.6 Main Result and Proof
Theorem 1.
Under the assumptions stated, if the vectors $\{\textbf{g}_{i}^{\top}\bm{\delta}_{j}\}_{j=1,2}$ are not pathologically structured (e.g., they are not both zero or strictly negative), then there exist merge coefficients $\alpha,\beta>0$ such that for at least one task $i\in\{1,2\}$,
@@ A.6 Main Result and Proof
Appendix B Experiment Details
@@
Our training protocol closely mirrors the public code of Ilharco et al. (2023). For each task, we fine-tuned CLIP backbones (ViT-B-32 and ViT-L-14) for 2000 updates using AdamW (Loshchilov & Hutter, 2019) with a weight-decay factor of 0.1. We adopted a cosine-annealed learning-rate schedule preceded by 200 warm-up steps and used a mini-batch size of 128; ViT-L-14 training employed gradient accumulation to match this effective batch size. Following the findings of Ilharco et al. (2023), we kept CLIP’s text encoder frozen and treated the logits obtained from class-specific prompts (e.g., “a photo of a {classname}”) as a fixed classification head, updating only the image encoder during fine-tuning. Regarding the learning rate, we used $10^{-4}$ only when training task vectors with large norms in the Norm Mismatch setting, and $10^{-5}$ for all other cases. In the Low Confidence setting, the label smoothing strength was set to $\alpha=0.1$.
@@
For all four merging methods adopted in this study, it is necessary to tune the task vector coefficient $\lambda_{t}$. Following Ilharco et al. (2023), we imposed a unified constraint on all $\lambda_{t}$ and searched the range from 0.0 to 1.0 (in increments such as 0.05) based on validation accuracy.
@@ B.1 Normalized Accuracy
Appendix C Additional Results
C.1 Norm Comparison across Layers
@@
C.2 Other Confidence Calibration Method and Merging Performance
@@
C.3 Impact of Task Vector Scaling on ViT-L-14
@@
Figure 7: Effect of scaling task vectors on test accuracy. For each of the eight vision tasks (ViT-L-14), we evaluate the model $\bm{\theta}_{\text{pre}}+\lambda\bm{\tau}$ as the scaling factor $\lambda$ varies from 0.00.00.0 to 3.03.03.0. Shrinking the task vector ($\lambda<1.0$) often preserves or even improves accuracy relative to the fine-tuned model ($\lambda=1.0$), while stretching the vector ($\lambda>1.0$) leads to sharp degradation. At $\lambda=3.0$, performance falls below that of the zero-shot model on all tasks. These results support shrinking long task vectors to match shorter ones when resolving norm disparities.
C.4 Scaling Alone Is Insufficient to Overcome Norm Mismatch

### v2

@@ 7 Conclusion
We presented DisTaC, a lightweight and practical pre-conditioning method for task vectors that improves the robustness of model merging in multi-task learning. Our analysis identified two major failure modes of norm disparity and low source-model confidence that frequently occur in real-world merging scenarios. DisTaC addresses both issues simultaneously via KD on unlabeled data, requiring only minimal computational cost and no access to task labels. Through extensive experiments, we demonstrated that DisTaC not only recovers performance degraded by task vector scaling, but also enhances confidence in the source models without sacrificing generalization. Furthermore, we showed that DisTaC enables state-of-the-art merging methods to succeed in challenging cases where they would otherwise fail. Our findings highlight the importance of task vector conditioning, and we believe that DisTaC provides a simple yet powerful tool to make model merging more reliable and broadly applicable.
8 Limitation
Our study is limited to CLIP-based vision tasks, and we have not evaluated DisTaC on other modalities or architectures. While we focus on the two main failure modes of norm disparity and low source-model confidence, we do not explore all possible causes of task interference. Additionally, DisTaC assumes access to unlabeled data for distillation, which can at times be challenging due to potential security constraints. Nevertheless, we emphasize that DisTaC achieves over 96% of ideal performance even when using extremely small datasets or data with severe distribution shifts, demonstrating strong robustness in such settings (see Appendix E.5). Furthermore, other approaches, such as yang2024adamerging; yan2025calm, also rely on the availability of unlabeled data. Despite these limitations, we believe that our experiments directly support our main claims on failure modes and are sufficient to demonstrate the effectiveness of our approach.
Acknowledgement
Our deepest gratitude goes out to the anonymous reviewers whose invaluable insights substantially enhanced the quality of this manuscript. This work was supported by RBC Borealis through the RBC Borealis AI Global Fellowship Award, which was awarded to Hiroki Naganuma. The computation resource of this project is supported by “TSUBAME Encouragement Program for Young/Female Users” of Center for Information Infrastructure at Institute of Science Tokyo and by “Joint Usage/Research Center for Interdisciplinary Large-scale Information Infrastructures” in Japan.
Appendix A Related Work
A.1 Model Merging and Task Arithmetic
Research on integrating multiple neural network models by performing operations on their parameters has been widely conducted, starting with utans1996weight. These techniques enable a model to learn diverse tasks with less time and computational resources, and have become increasingly important in recent years as the number of model parameters has grown dramatically. For instance, in early approaches to model merging, models with the same architecture were fine-tuned and then merged by averaging their parameters (wortsman2022soups; choshen2022fusing). More sophisticated methods have since been proposed, such as Fisher Merging (matena2022merging), which is based on maximizing the posterior probability of the model, and RegMean (jin2023dataless), which minimizes the distance between output activations before and after merging. In contrast, task arithmetic (ilharco2023editingmodelstaskarithmetic) focuses on the task vector, defined as the difference in parameters between a fine-tuned model and a pre-trained model, and performs addition and subtraction of task vectors in parameter space. This approach offers the advantage of allowing flexible, localized modifications to the model and has found applications across diverse tasks (tang2024merging; su-etal-2024-task; yoshikawa2025transferringvisualexplainabilityselfexplaining; naganuma2025fairnesstaskarithmeticrole).
Recent research on task arithmetic has theoretically analyzed the simple addition of task vectors and proposed multiple methods to address its shortcomings. Approaches aimed at improving the properties of task vectors focus on the linearity in fine-tuning (ortiz-jimenez2023task; yoshida2025mastering). These methods, based on the Neural Tangent Kernel (NTK) (jacot2018neural), treat the model’s output as linear during fine-tuning in order to reflect vector operations in parameter space onto the model’s inputs and outputs. Meanwhile, several studies have been conducted from the perspective of mitigating interference between task vectors. TIES-Merging (yadav2023tiesmerging) emphasizes the removal of redundant elements and the consideration of sign in each vector dimension. AdaMerging (yang2024adamerging), on the other hand, automatically adjusts merging coefficients per task and per layer to reduce task interference and enhance robustness through test-time adaptation. wang2024localizing introduced a framework for pinpointing the parameters that carry information shared across tasks and, on that basis, proposed Consensus Merging, which builds task-wise masks that align more closely with inter-task consensus than the masks used in TIES-Merging. While traditional multi-objective optimization can be computationally prohibitive, li2025map amortized this cost by leveraging quadratic approximations to identify diverse Pareto-optimal merging solutions. More recently, wei2025modeling reformulated model merging as minimizing the loss gap between the merged model and each task-specific model, introducing DOGE with subspace projection and task-aware scaling. TSVM (gargiulo2025task) interprets task interference as non-orthogonality among the layer-wise singular vectors of the task vectors; by whitening those singular directions, TSVM further improves merge quality.
Despite these advances, ilharco2023editingmodelstaskarithmetic and nearly all follow-up studies on multi-task model merging benchmark their methods under highly idealized settings, leaving real-world failure modes largely unexplored. In this work, we show that (i) discrepancies in task vector norms and (ii) low source-model confidence are key sources of interference. We introduce DisTaC as a simple pre-conditioning step that mitigates both problems before merging.
A.2 Knowledge Distillation
DisTaC addresses the limitations of existing task arithmetic methods by incorporating knowledge distillation. Knowledge distillation is a technique proposed for transferring knowledge from a teacher model to a smaller student model (hinton2015distilling). Although initially intended for model compression (hinton2015distilling; kim2018paraphrasing; sanh2020distilbert), it has also been applied in contexts such as self-distillation, where repeated distillation between models of the same architecture leads to performance improvement (furlanello2018born; zhang2019beyourown; zhan2020selfdistillation). Among these applications, several studies have explored generating models that can handle multiple tasks by distilling knowledge from single or multiple teacher models (luo2019knowledge; hao2023oneforall; xu2023multi). These approaches achieve distillation by mapping the parameters of multiple teacher models into a shared space for the student model. Conversely, it is also possible to distill models with different architectures individually to obtain task vectors, which can then be merged using task arithmetic (merugu2025statsmerging). DisTaC adopts the latter approach and resolves the issue of variability in the norms of task vectors by obtaining them through distillation.
Applying distillation to task arithmetic requires addressing the impact of soft targets. Numerous studies have analyzed the effects of label smoothing in the context of knowledge distillation (muller2019does; shen2021is; chandrasegaran2022revisiting; zheng2024knowledge). In this study, we demonstrate that fine-tuning with soft targets significantly affects the models obtained through model merging, and propose a method to mitigate this effect by increasing the confidence of the student model.
Appendix B Proof for Proposition 1
Let $\delta=\|\bm{\tau}_{1}\|/\|\bm{\tau}_{2}\|$ and assume $\bm{\tau}_{1}\perp\bm{\tau}_{2}$. Then
$\|\bm{\tau}_{\mathrm{merge}}\|^{2}=\|\bm{\tau}_{1}+\bm{\tau}_{2}\|^{2}=\|\bm{\tau}_{1}\|^{2}+\|\bm{\tau}_{2}\|^{2}=(1+\delta^{2})\|\bm{\tau}_{2}\|^{2}.$
For the cosine similarity with $\bm{\tau}_{2}$, we compute
$\cos(\bm{\tau}_{\mathrm{merge}},\bm{\tau}_{2})=\frac{\bm{\tau}_{\mathrm{merge}}\cdot\bm{\tau}_{2}}{\|\bm{\tau}_{\mathrm{merge}}\|\|\bm{\tau}_{2}\|}=\frac{\|\bm{\tau}_{2}\|^{2}}{\sqrt{(1+\delta^{2})}\,\|\bm{\tau}_{2}\|^{2}}=\frac{1}{\sqrt{1+\delta^{2}}}.$
Using the inequality $(1+\delta^{2})^{-1/2}\geq 1-\tfrac{1}{2}\delta^{2}$ for $\delta\geq 0$, we obtain the lower bound.
Similarly, for the cosine similarity with $\bm{\tau}_{1}$,
$\cos(\bm{\tau}_{\mathrm{merge}},\bm{\tau}_{1})=\frac{\bm{\tau}_{\mathrm{merge}}\cdot\bm{\tau}_{1}}{\|\bm{\tau}_{\mathrm{merge}}\|\|\bm{\tau}_{1}\|}=\frac{\|\bm{\tau}_{1}\|^{2}}{\sqrt{(1+\delta^{2})}\,\|\bm{\tau}_{1}\|\|\bm{\tau}_{2}\|}=\frac{\delta}{\sqrt{1+\delta^{2}}}.$
Since $\delta/\sqrt{1+\delta^{2}}\leq\delta$, the claim follows.
Hence, when $\delta\ll 1$, the merged vector is nearly aligned with $\bm{\tau}_{2}$ while its alignment with $\bm{\tau}_{1}$ is suppressed by a factor of $O(\delta)$. ∎
@@ Appendix B Proof for Proposition 1
Appendix C Theoretical Insights into Task Vector Merging for Models Optimized with Distinct Objectives
@@
C.1 Notation and Assumptions
We use the following notation. For task $i$, the standard cross-entropy (CE) objective is
$J_{i}^{\mathrm{CE}}(\bm{\theta}):=-\,\mathbb{E}_{(x,y)\sim\mathcal{D}_{i}}\!\left[\log p_{\bm{\theta}}(\bm{y}\mid\bm{x})\right].$
We also consider a calibrated objective that augments CE with a generic penalty $\mathcal{C}_{i}(\bm{\theta})$ [^1] For example, a detailed description of evaluating focal loss can be found in kimura2025geometric. weighted by $\lambda_{i}>0$:
@@ C.1 Notation and Assumptions
For either objective $\star\in\{\mathrm{CE},\mathrm{CAL}\}$, the task-specific optimum is denoted
$\bm{\theta}_{i}^{\star}:=\arg\min_{\bm{\theta}}J_{i}^{\star}(\bm{\theta}).$
Throughout, we assume the objectives $J_{i}^{\mathrm{CE}}$ and $J_{i}^{\mathrm{CAL}}$ are $C^{2}$ in a neighborhood of a fixed base parameter $\bm{\theta}_{0}$. Let $H_{i}:=\nabla^{2}J_{i}^{\mathrm{CE}}(\bm{\theta}_{0})$ denote the CE Hessian at $\bm{\theta}_{0}$ and assume $H_{i}$ is positive-definite, ensuring that $\bm{\theta}_{0}$ lies in a locally convex region of the CE landscape. For notational convenience we write the gradients at $\bm{\theta}_{0}$ as
$\mathbf{g}_{i}:=\nabla J_{i}^{\mathrm{CE}}(\bm{\theta}_{0}),\qquad\mathbf{b}_{i}:=\nabla\mathcal{C}_{i}(\bm{\theta}_{0}).$
C.2 Theoretical Preliminaries for the Main Result
C.2.1 Derivation of the Standard Task Vector
@@
C.2.2 Derivation of the Calibrated Task Vector
@@
Gradient and hessian at the base point.
@@
Neumann series expansion of $\tilde{H}_{i}^{-1}$.
@@
The inverse is then given by $\tilde{H}_{i}^{-1}=(I+\lambda_{i}H_{i}^{-1}A_{i})^{-1}H_{i}^{-1}$. We can expand the term $(I+\lambda_{i}H_{i}^{-1}A_{i})^{-1}$ using a Neumann series (horn2012matrix), $(I+X)^{-1}=\sum_{k=0}^{\infty}(-X)^{k}$, which converges if the spectral radius of $X$ is less than 1. Assuming $\lambda_{i}$ is sufficiently small such that $\|\lambda_{i}H_{i}^{-1}A_{i}\|<1$, we have:
@@
In many practical scenarios, especially after extensive pre-training, the initial gradient norm $\|\textbf{g}_{i}\|$ is small. Consequently, the term $\lambda_{i}H_{i}^{-1}A_{i}H_{i}^{-1}\textbf{g}_{i}$, which is of order $\mathcal{O}(\lambda_{i}\|\textbf{g}_{i}\|)$, is often negligible compared to the term $\lambda_{i}H_{i}^{-1}\textbf{b}_{i}$, which is $\mathcal{O}(\lambda_{i})$. Under this simplifying assumption, we can define the first-order correction due to calibration as:
@@
C.2.3 Task Vector Merging
@@
Taylor expansion of the CE loss for merged vectors.
@@
The term for task $i$ itself ($i=1$ and analyzing $\bm{\tau}_{1}^{\mathrm{CE}}$, or $i=2$ and analyzing $\bm{\tau}_{2}^{\mathrm{CE}}$) is $\alpha(\bm{\tau}_{i}^{\mathrm{CE}})^{\top}H_{i}\bm{\tau}_{i}^{\mathrm{CE}}=-\alpha\|\bm{\tau}_{i}^{\mathrm{CE}}\|_{H_{i}}^{2}$. Since $H_{i}$ is positive-definite, this self-term is strictly negative. The cross-term’s sign is indefinite. However, the dominant contribution to the loss change is typically negative and of order $\mathcal{O}(\|\bm{\tau}\|^{2})$, indicating that merging CE vectors does not increase the loss at first order.
Merging of calibrated vectors.
@@
C.3 Main Result and Proof
Proposition 2.
Under the assumptions stated, if the vectors $\{\textbf{g}_{i}^{\top}\bm{\delta}_{j}\}_{j=1,2}$ are not both zero or strictly negative, then there exist merge coefficients $\alpha,\beta>0$ such that for at least one task $i\in\{1,2\}$,
@@ C.3 Main Result and Proof
Appendix D Experiment Details
@@
Our training protocol closely mirrors the public code of ilharco2023editingmodelstaskarithmetic. For each task, we fine-tuned CLIP backbones (ViT-B-32 and ViT-L-14) for 2000 updates using AdamW (loshchilov2018decoupled) with a weight-decay factor of 0.1. We adopted a cosine-annealed learning-rate schedule preceded by 200 warm-up steps and used a mini-batch size of 128; ViT-L-14 training employed gradient accumulation to match this effective batch size. Following the findings of ilharco2023editingmodelstaskarithmetic, we kept CLIP’s text encoder frozen and treated the logits obtained from class-specific prompts (e.g., “a photo of a {classname}”) as a fixed classification head, updating only the image encoder during fine-tuning. Regarding the learning rate, we used $10^{-4}$ only when training task vectors with large norms in the Norm Mismatch setting, and $10^{-5}$ for all other cases. In the Low Confidence setting, the label smoothing strength was set to $\alpha=0.1$.
@@
For all four merging methods adopted in this study, it is necessary to tune the task vector coefficient $\lambda_{t}$. Following ilharco2023editingmodelstaskarithmetic, we imposed a unified constraint on all $\lambda_{t}$ and searched the range from 0.0 to 1.0 (in increments such as 0.05) based on validation accuracy.
@@ D.1 Normalized Accuracy
Appendix E Additional Results
E.1 Norm Comparison across Layers
@@
E.2 Other Confidence Calibration Method and Merging Performance
@@
E.3 Impact of Task Vector Scaling on ViT-L-14
@@
Figure 7: Effect of scaling task vectors on test accuracy. For each of the eight vision tasks (ViT-L-14), we evaluate the model $\bm{\theta}_{\text{pre}}+\lambda\bm{\tau}$ as the scaling factor $\lambda$ varies from $0.0$ to $3.0$. Shrinking the task vector ($\lambda<1.0$) often preserves or even improves accuracy relative to the fine-tuned model ($\lambda=1.0$), while stretching the vector ($\lambda>1.0$) leads to sharp degradation. At $\lambda=3.0$, performance falls below that of the zero-shot model on all tasks. These results support shrinking long task vectors to match shorter ones when resolving norm disparities.
E.4 Scaling Alone Is Insufficient to Overcome Norm Mismatch